A player caught a cricket ball of mass 150 g moving at a rate of 20 m/s. If the catching process is completed in 0.1 s, the force of the blow exerted by the ball on the hand of the player is equal to ?

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
From the given data,
The force of the blow exerted by the ball on the hand of the player is
(mv-0)=0.15*20=3
F=3/0.1=30N
Thanks and Regards
Shaik Aasif
askIITians faculty